OpenStudy (anonymous):

Which expression is a factor of 3xy + 2x – 18y – 12?

OpenStudy (anonymous):

where are the expression's?

OpenStudy (anonymous):

x+3 x+6 3x-2y 2+3y

OpenStudy (cwrw238):

2 + 3y

OpenStudy (cwrw238):

3xy + 2x - 18y - 12 = x (3y + 2) - 6(3y + 2)

OpenStudy (cwrw238):

3y + 2 is of course = 2 + 3y

OpenStudy (anonymous):

Thanks

OpenStudy (cwrw238):

that one gave me a headache

OpenStudy (cwrw238):

yw